OpenStudy (daniel.ohearn1):

Consider the set S={w,x,y,z}. If the set of all possible sums of any three different elements from S is {-1,3,5,8}, then what is the set S?

OpenStudy (zzr0ck3r):

start adding 3 until you get 4 -1+3+5 = 7 -1+3+8= 10 5+3+8=16 8+5-1=12

OpenStudy (daniel.ohearn1):

w+x+y, w+x+z, w+y+z x+y+z -> 3w + 3x+3y+3z=15 -> w+x+y+z=5